OpenStudy (anonymous):

What is a Taylor Series for f(x)=ln(2+2x) ?

OpenStudy (anonymous):

ummm I might be able to do this... we'll need to collaborate though

OpenStudy (anonymous):

they gave the first part: ln(2)+\[\sum_{n=1}^{\infty}\]

OpenStudy (anonymous):

so whats in the sigma, thats important

OpenStudy (anonymous):

thats what we have to try to find lol

OpenStudy (anonymous):

does the question say exactly that? What is a taylor series for...

OpenStudy (anonymous):

take the derivative of ln(x) like 4 or 5 times see if you can see a pattern

OpenStudy (anonymous):

yes, What is a Taylor Series for f(x)=ln(2+2x)

OpenStudy (anonymous):

of and at x=1/2 we know that ln(1) = 0

OpenStudy (anonymous):

\[\sum_{?}^{?} (f^{(n)}(1/2))/n! * (2x-(1/2))^{n}\]

OpenStudy (anonymous):

the general formula would be F^(n) (x) = (-1)^(n+1) (n-1) ! --------------- x^(n)

OpenStudy (anonymous):

thats my best guess but I think I'm taking it too fat by kinda approxximating

OpenStudy (anonymous):

yes it would now try writing that in sigma notation which is what I did up there...not sure if this is right to do...

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

ha, I have no idea

OpenStudy (anonymous):

Go to the general form of the Taylor Series. T(x) = f(a) + f'(a)(x-a) + f''(a)(x-a)^2/2! + ... and plug the derivatives of ln(2+2x) into the places you see fit (I think in this case it's safe to set a=0).

OpenStudy (anonymous):

I dont know where I would stop using the general form...?

OpenStudy (anonymous):

Until you see the general form that they're taking...and then you can put the general form of the series into your sum symbol.